# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v2

@@ II-B Radar system
The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification. Its flexibility in varying emitted beams allows for the simultaneous detection of more targets, to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems.
Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [32, 33]. The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34]. The total radar return signal at the $l$-th snapshot is shown as follows,
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},$ (5)
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix of the BS-Target-BS route for the $t$-th target, following a Rayleigh distribution [35]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot. To analyse the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target. The radar precoding matrix $\mathbf{W}_{r,l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS, under the binary hypotheses $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ signifies target presence and $\mathcal{H}_{0}$ denotes target absence, can be mathematically represented as follows,
@@
The $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
III The communication KLD: ZF and MRT
In this section, we derive the KLD of the communication system for the MRT and ZF beamforming using the shared deployment antenna configuration. The analysis is carried out for the normalised precoders introduced in Section II-A.
To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
@@ III The communication KLD: ZF and MRT
III-A MRT beamforming
For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components. The examination starts with the IUI term, which is the second term in (2), given by,
$\omega_{\mathrm{MN},k}=d_{\mathrm{c},k}^{-\zeta/2}\mathbf{h}_{k}^{T}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}=\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\|\mathbf{H}\|_{F}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (8)
Since the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d. complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors. Thus, the central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a large number of antennas $N$. Meanwhile, the normalisation term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$. Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable. This ratio distribution does not have a simple closed-form expression. However, we can approximate its behaviour for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$. Using this approximation, we can write (8) as,
$\omega_{\mathrm{MN},k}\approx\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\sqrt{NK\sigma_{h}^{2}}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (9)
$\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking CLT, for large $KN$, the sum approaches a complex normal distribution with zero mean and a variance of
$\sigma_{\omega,\mathrm{MN}}^{2}\approx\frac{d_{\mathrm{c},k}^{-\zeta}}{NK\sigma_{h}^{2}}\sum_{i\neq k}N\sigma_{h}^{4}p_{i}=\frac{d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}}{K}\sum_{i\neq k}p_{i},$ (10)
The accuracy of the approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation, which compares it with simulated density functions under different values of $N$. It can also be shown that the distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
@@ III-A MRT beamforming
where $\mathbf{R}_{w,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{\tilde{w}}_{\mathrm{r},t,l}\mathbf{\tilde{w}}_{\mathrm{r},t,l}^{H}$. Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$. In order to find the KLD for the MRT case with matrix normalisation, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2). This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
$|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}=|\mathbf{h}_{k}^{T}\mathbf{h}_{k}^{*}|^{2}/\|\mathbf{H}\|_{F}^{2}=\|\mathbf{h}_{k}\|^{4}/\|\mathbf{H}\|_{F}^{2}.$ (12)
The expected value of this term can be approximated using the properties of the chi-squared distribution and the approximation $\mathbb{E}[1/X]\approx 1/\mathbb{E}[X]$ for a random variable $X$,
$\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]\approx p_{k}\mathbb{E}[\|\mathbf{h}_{k}\|^{4}]/\mathbb{E}[\|\mathbf{H}\|_{F}^{2}]=p_{k}(N+1)\sigma_{h}^{2}/K.$ (13)
Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, substituting these into equation (7), then the KLD can be found as [21, Corollary 1],
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]}{(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2},$ (14)
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$. Substituting (13) in (14), the KLD for MRT precoding with matrix normalisation is,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N+1)\sigma_{h}^{2}}{K\>(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2}.$ (15)
III-B ZF beamforming
With ZF precoding, the received signal at the $k$-th user is,
$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l},$ (16)
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix. Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
@@ III-B ZF beamforming
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[|\mathbf{\tilde{W}}_{\mathrm{c}}|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (17)
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (17), and (18) in (7), then the KLD becomes,
@@
The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
@@
III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
@@
Note that, the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$ (24)
Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
×$\times\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ and $\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}=\lambda/(M(M-1))$, substituting it in the KLD expression which becomes,
$\begin{split}\mathrm{KLD}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (26)
It should be noted that the second term in this equation, represents the form of the SINR, where the interference is the sum of IUI and radar interference.

### v4

@@ II-B Radar system
The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification. Its flexibility in varying emitted beams allows for the simultaneous detection of more targets, to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems. Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [35, 36]  [^1] The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [37].. The total radar return signal at the $l$-th snapshot is shown as follows,
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vspace{-0.05 in}$ (4)
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix that captures the complete BS-Target-BS path characteristics for the $t$-th target, inherently incorporating propagation effects, target cross-section, and reflection coefficients. The target response matrix follows a Rayleigh distribution to model the composite effects of multiple scattering centres inherent in extended targets, thereby characterising the statistical nature of target scattering while preserving the direct path essential for target detection [38, 39]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot. To process the radar returns from all beams, a bank of matched filters is typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Assuming $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$ using signal space and time domain, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target [39, 34]. The radar precoding matrix $\mathbf{W}_{\mathrm{r},l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS after match-filtering, under the binary hypothesis testing problem $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ represents the target presence whereas $\mathcal{H}_{0}$ denotes the target absence, is denoted by $\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}$ and can be mathematically represented as follows,
@@
where $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot at the output of the matched filter.
III The communication KLD: ZF and Arbitrary precoding matrix
In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver. The analysis is carried out for the normalised precoders introduced in Section II-A. To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
@@ III The communication KLD: ZF and Arbitrary precoding matrix
III-A ZF beamforming
With ZF precoding, the received signal at the $k$-th user is,
$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l},$ (7)
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix. Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
@@ III-A ZF beamforming
where $\mathbf{R}_{w,t}\!\!=\!\!\frac{1}{L}\!\sum_{l=1}^{L}\!\!\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}^{H}$. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
@@
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (9)
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (9), and (10) in (6), and considering all possible pairs of dissimilar symbols [22, Corollary 1], then the KLD becomes,
@@
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is a constant that depends on the constellation. The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
@@
III-B Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
@@
Note that, the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},$ (16)
by substituting the evaluated mean vectors and covariance matrices in (14), and (16) into the KLD expression given in (6), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, which can be represented as follows,
$\!\!\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}\!=\!\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
×$\!\!\times\!\Bigg{(}\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg{)}^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg{)}\!.\!\!$ (17)
Thereafter, by considering all possible pairs of dissimilar symbols, then the KLD expression using $\lambda$ is simplified as,
$\begin{split}\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (18)
It should be noted that the second term in this equation, represents a form of the SINR, where the interference is the sum of IUI and radar interference.